# KA-TP-05-2025

### v1

@@ 1 Introduction
The true nature of Dark Matter (DM) is one of the most prominent open questions of contemporary particle physics. While the Standard Model (SM) of particle physics has been structurally completed with the discovery of the Higgs boson [1, 2], which turned out to behave very SM-like, and has been tested to the highest accuracy, it does not provide a viable DM candidate. Numerous astrophysical and cosmological observations point to the existence of DM [3, 4, 5, 6, 7, 8]. However, it remains totally unclear what is the nature of DM. It can range from particle character to supermassive objects like e.g. primordial black holes [9], and thus cover a huge mass range. Assuming particle character for DM, there are two main mechanisms that can produce DM in agreement with all observables, in particular the measured relic density [10]. These are given by freeze-in [11] and freeze-out [12, 13], where we consider here the latter process. In the freeze-out mechanism, weakly interacting massive particles (WIMPs) are assumed to be in thermal equilibrium with the thermal bath until the expansion rate of the universe becomes larger than the DM (co-)annihilation rate into lighter particles. The thermal relic density that we observe today is then obtained from the solution of the Boltzmann equation for the evolution of the DM number density.
@@
There exist numerous codes that calculate for beyond-SM (BSM) extensions DM observables, in particular the DM relic density, such as SuperIso Relic [14], DarkSUSY [15], MicrOMEGAs [16, 17, 18, 19], or MadDM [20, 21, 22]. [^6] There are many more DM tools publicly available, cf. e.g. the review [23]. These tools calculate the relic density obtained from the freeze-out of WIMPs. The codes SuperIso Relic and DarkSUSY have (originally) been written for supersymmetric (SUSY) extensions. In the case of DarkSUSY, the users can also directly provide the thermally averaged cross section (TAC) for their model to compute the relic density. In MicrOMEGAs and MadDM all matrix elements are automatically calculated during the execution of the program by CalcHEP [24] and MadGraph5_aMC@NLO [25], respectively. They can also handle multi-component DM. All these codes rely on leading-order matrix elements. In MicrOMEGAs, the TACs can be improved to include next-to-leading order (NLO) terms by editing predefined functions. For DM in the minimal supersymmetric extension (MSSM), the code DM@NLO [26, 27] exists, which provides the NLO (SUSY) QCD calculation to DM (co-)annihilation. The programs furthermore compute the direct detection cross sections. For further details and codes including indirect DM searches, we refer to [23].
In this paper, we present a new tool, the C++ code RelExt. The code computes the relic density based on the freeze-out mechanism for arbitrary models featuring a DM candidate stabilized by a discrete $\mathbb{Z}_{2}$ symmetry. The user solely has to provide the FeynRules [28] model files, and the relic density is then automatically computed by RelExt. In this first release, the relic density is computed to leading-order in the thermally averaged cross section of the DM (co-)annihilation channels. The distinctly new feature of the code w.r.t. existing codes is the scan procedure. The code implements an efficient scan algorithm that searches for parameter configurations that generate the correct relic density within the experimentally allowed limits, provided the model under investigation is able to saturate the relic density. The code thereby provides a powerful tool to further constrain DM models by requiring the reproduction of the correct DM relic density. This can be used to subsequently investigate the phenomenological implications of this constraint for current and future collider and DM experiments on the one hand. On the other hand, this kind of investigation gives guidelines for further model building, in case the favored DM fails to reproduce the measured DM relic density.
@@
  • Having written the code ourselves, allows us to consistently combine it with the other codes generated by our collaboration to test new physics models with respect to particle, astroparticle and cosmological constraints, in particular our scanning tool ScannerS [29, 30], which checks for theoretical and experimental constraints of extended Higgs sector models, and the code BSMPT [31, 32, 33], which investigates BSM Higgs sectors with respect to their phase history in the early universe and possibly related gravitational waves signals.
At present, the following beyond-SM models featuring DM candidates have been pre-implemented and tested: the complex singlet extension of the Standard Model (CxSM) [29, 34, 35, 36, 37, 38, 39, 40], the Next-to-2-Higgs-Doublet Model (N2HDM) in its dark doublet phase (DDP) [41, 42], the model CP in the Dark (CPVDM) [43, 44, 45], the Two-Real-scalar-Singlet Model (TRSM) [46, 47], and the BDM5 [48, 49]. New models, as stated above, can be included by providing the corresponding FeynRules model files. Note that the code at present does not handle models with more than one DM candidate. Furthermore, models where DM annihilation processes and the corresponding back-reactions differ (due to CP violation e.g.) are not supported either. Such extensions will be included in future releases.
@@ 2 Code Overview
This first part of the code allows the user to implement new models. This is done by providing the FeynRules [28] files for the model to be implemented. The code uses the packages FeynArts [50] and FeynCalc [51, 52] to automatically generate all tree-level DM annihilation amplitudes relevant for the calculation of the DM relic density. In this first release of the code these are the squared amplitudes for all tree-level DM $2\to 2$ (co-)annihilation processes. These amplitudes will be saved together with additional necessary information about the model in C++ files, which can then be compiled and used by the second part of the code. In case no new model is provided, the second part of the code can still be used with the already implemented models that are shipped with the code. The present version of the code is able to handle models with one $\mathbb{Z}_{2}$ symmetry, i.e. with one DM candidate.
@@
The second part of the code does not require Mathematica [53] and is purely written in C++. It performs the scans in the parameter spaces of the newly generated or already implemented models which search for parameter configurations that saturate the experimental limits on the relic density or that at least do not exceed it, in case weaker conditions are required. For this, the necessary thermally averaged cross sections are computed. This comprises also the computation of the needed total widths required for the virtual particle exchanges, and of the running couplings entering the interaction vertices of the scattering amplitudes. The code is designed such that it can be run with different settings and functions via a main.cpp file found within each model folder. The main.cpp allows the user to perform searches in the parameter space of a model using the provided functions (for more details see Sec. 3.4 and App. A). The values of the calculated relic densities and the corresponding relevant input parameters are saved in the output file specified by the user.
@@ 3.1 System Requirements
  • Mathematica v12[53] or higher is required for the implementation of new models. For the usage of implemented models, Mathematica and the packages listed here below are optional.
  • The Mathematica package FeynRules [28] is used to compute the Feynman rules for new models and to generate the required input files, which will be used by our code to implement new models.
  • FeynArts v3.11 [50] and FeynCalc v9.3.1 [51, 52] are used to calculate the squared amplitudes of the DM annihilation processes.
@@ 3.2 Download
{labeling}
@@
This directory contains the header files of the code.
@@
This directory contains the already implemented model CPVDM [43, 45].
This directory contains the already implemented model CxSM [29, 34, 35, 36, 37, 38, 39, 40].
This directory contains the already implemented model N2HDM in its dark doublet phase [41, 42].
This directory contains the already implemented model TRSM [46, 47].
This directory contains the already implemented model BDM5 [48, 49].
This directory contains the numerical data for the computation of the effective degrees of freedom [54].
@@ 4 How to Implement a new Model
    The first input parameter is the path to the FeynRules directory. The second one is the name of the FeynRules model file, while the last input is the name of the Lagrangian variable, which is defined in the model file. This command will automatically search the model file in \$RelExt and generate all relevant (co-)annihilation amplitudes for DM freeze-out. The generated md_[ModelName] directory contains the following files and folders: {labeling}FR_modfiles
  3. FR_modfiles
@@
[…]
@@ 5 Structure and Description of the Algorithms
The BSM models that provide DM candidates depend on a set of input parameters that can be more or less lengthy. Depending on the model and the interplay of the parameter values and the constraints on the models, more or less extensive parameter scans have to be performed, in order to find viable relic density values and to perform meaningful phenomenological investigations. A main objective of the code is therefore to solve the Boltzmann equation given in terms of the yield $Y$ by [55]
@@
Here, $\lambda$ is the Källén function, $\theta$ the azimuthal scattering angle, and $\mathcal{M}_{ij\rightarrow kl}$ the $2\rightarrow 2$ (co-)annihilation matrix element. After solving Eq. (5.1) from $x=0$ to $x=x_{0}$, the relic density can be calculated using
@@ 5.1 Amplitude Generation
where $m_{j}$ and $\Gamma_{j}$ are the mass and total decay width, respectively, of the mediator particle $j$. The code automatically identifies all $s$-channel mediator particles and calculates their total decay widths. We will give an overview on how the decay widths are obtained in Sec. 5.2.
@@ 5.2 Computation of the Total Widths
In the computation of the total widths for scalar and pseudoscalar mediator particles, we include the available dominant higher-order corrections to the decay widths, as explained in more detail in the following. First of all, for the scalar/pseudoscalar mediator particles, analytical formulae are implemented in RelExt and used for the calculation of the partial decay widths, which are summed up to the total widths. We follow here mostly the implementation of HDECAY [56, 57], a code for the computation of the Higgs decays for the SM and the minimal supersymmetric SM extension (MSSM), including state-of-the-art higher-order corrections, and take the following formulae given in Ref. [58], unless stated otherwise. Here, we include the coupling modification factors of the scalar/pseudoscalar mediator to the SM particles according to the model under investigation. For the decays into lepton final states, we take the tree-level formula given in Eq. (21). Since the QCD interactions are not affected by modifications of the Higgs sector, we include higher-order QCD corrections in the decays into quarks. We interpolate here between the large Higgs mass region and the threshold region, and we include the relative QCD corrections $\delta_{\text{QCD}}$ and the relative top-quark induced contributions $\delta_{t}$, as given in Eq. (23). For the large Higgs mass region, we apply formula Eq. (15) for $\delta_{\text{QCD}}$, using running $\overline{\mbox{MS}}$ quark masses both at tree-level and at higher order. These and the running strong coupling constant $\alpha_{s}$ with five active flavours are taken at the scale of the mass of the decaying Higgs, thereby absorbing large logarithms. The $\delta_{t}$ corrections differ for the scalar and pseudoscalar case and are given after Eq. (23). In the threshold region, finite quark mass effects become relevant, and we apply Eq. (25). The QCD corrections are given in Eq. (17) for the scalar and in Eq. (26) for the pseudoscalar case, where we now take quark pole masses everywhere, both at leading and at higher order. The coupling modification factor for the Higgs coupling to the quarks in Eq. (23) and Eq. (25), respectively, is automatically replaced by the corresponding one for the model under consideration.
@@
Additionally, off-shell decays into massive gauge bosons $V$ ($V=W^{\pm},Z$) are taken into account. In a CP-conserving theory, they are only allowed for the scalar mediators. For energies $\sqrt{s}\geq 2m_{V}$, we apply the on-shell decay formula, given by Eqs. (34–36). Since the QCD corrections are of moderate size, we do not include them in our calculation. This also ensures a smooth transition to the off-shell decays, where no QCD corrections are available in a compact form. For energies below the production of any on-shell gauge bosons, $\sqrt{s}\leq m_{V}$, we apply Eqs. (37–39) for the production of two off-shell gauge bosons. In the energy range $m_{V}\leq\sqrt{s}\leq 2m_{V}$, we compute the decay into one off-shell and one on-shell gauge boson, using Eqs. (51,52), given in [59].
@@
In BSM models, further Higgs decays are possible, which, depending on the parameter region, can become sizeable, or even dominant and therefore have to be included in the total widths as well. These can be e.g. decays into a lighter Higgs plus gauge bosons, lighter Higgs pairs, new heavy fermions or new heavy gauge boson pairs. These non-SM-like decays are computed automatically by the code and added to the decay widths into SM particles to obtain the total width.
@@ 5.3 Computation of the Thermally Averaged Cross Section
With the generated amplitudes we move on to the computation of the TAC given in Eq. (5.4). We start by integrating Eq. (5.6) over $\cos\theta$. For this, we compute a first estimate of the integral using a composite Simpson’s $3/8$ rule [60]. We use this estimate as part of the termination criterion for the adaptive Simpson’s $3/8$ method with which we compute the cross section. The use of the estimate in the termination criterion ensures a rapid convergence in areas where the integrand is small. Next, we need to integrate the numerator of Eq. (5.4). In order to do this we first sort the (co-)annihilation processes by the magnitude of their rest masses via
@@
This allows us to adjust the lower integration bound accordingly for each process. Additionally, we neglect channels which are heavily Boltzmann suppressed, i.e., that do not fulfill the criterion [18]
@@
with the recommended value $B_{\epsilon}=10^{-6}$. Further, we approximate the modified Bessel functions of the second kind via [61] [^8] To ensure that the precision of our results is not influenced by this, we use this approximation only for $x>5$.
$K_{n}(x)\approx\sqrt{\frac{\pi}{2x}}e^{-x}\left(1+\frac{4n^{2}-1}{1!(8x)}+\frac{\left(4n^{2}-1\right)\left(4n^{2}-9\right)}{2!(8x)^{2}}+...\right)\,.$ (5.11)
This approximation results in an increase of the numerical stability of the TAC for large values of $x$ and decreases the computational effort. Similar to the $\cos\theta$ integration, we compute a first estimate of the TAC integral using the Kronrod 61 point method [60] for the regions outside of $s$-channel resonance peaks. As in the $\cos\theta$ integration, we use this estimate in the termination criterion to compute the integral using an adaptive Gauss-Kronrod 15 point method [60]. Since resonance peaks are challenging to deal with numerically, we integrate them separately if they are close enough to the energy of the rest mass of the considered channel, i.e. if the condition

### v2

@@ 1 Introduction
The true nature of Dark Matter (DM) is one of the most prominent open questions of contemporary particle physics. While the Standard Model (SM) of particle physics has been structurally completed with the discovery of the Higgs boson [Aad:2012tfa, Chatrchyan:2012ufa], which turned out to behave very SM-like, and has been tested to the highest accuracy, it does not provide a viable DM candidate. Numerous astrophysical and cosmological observations point to the existence of DM [hoekstra2002nf, koopmans2002qh, metcalf2003sz, moustakas2002iz, rubin1970rotation, clowe2004weak]. However, it remains totally unclear what is the nature of DM. It can range from particle character to supermassive objects like e.g. primordial black holes [Chapline:1975ojl], and thus cover a huge mass range. Assuming particle character for DM, there are two main mechanisms that can produce DM in agreement with all observables, in particular the measured relic density [Planck:2018vyg]. These are given by freeze-in [Hall:2009bx] and freeze-out [Zeldovich:1965gev, Bertone:2004pz], where we consider here the latter process. In the freeze-out mechanism, weakly interacting massive particles (WIMPs) are assumed to be in thermal equilibrium with the thermal bath until the expansion rate of the universe becomes larger than the DM (co-)annihilation rate into lighter particles. The thermal relic density that we observe today is then obtained from the solution of the Boltzmann equation for the evolution of the DM number density.
@@
There exist numerous codes that calculate for beyond-SM (BSM) extensions DM observables, in particular the DM relic density, such as SuperIso Relic [Arbey:2018msw], DarkSUSY [Bringmann:2018lay], MicrOMEGAs [Belanger:2001fz, Belanger:2004yn, Belanger:2006is, Alguero:2023zol], or MadDM [Backovic:2013dpa, Backovic:2015cra, Ambrogi:2018jqj]. [^6] There are many more DM tools publicly available, cf. e.g. the review [Arina:2020alg]. These tools calculate the relic density obtained from the freeze-out of WIMPs. The codes SuperIso Relic and DarkSUSY have (originally) been written for supersymmetric (SUSY) extensions. In the case of DarkSUSY, the users can also directly provide the thermally averaged cross section (TAC) for their model to compute the relic density. In MicrOMEGAs and MadDM all matrix elements are automatically calculated during the execution of the program by CalcHEP [Belyaev:2012qa] and MadGraph5_aMC@NLO [Alwall:2011uj], respectively. They can also handle multi-component DM. All these codes rely on leading-order matrix elements. In MicrOMEGAs, the TACs can be improved to include next-to-leading order (NLO) terms by editing predefined functions. For DM in the minimal supersymmetric extension (MSSM), the code DM@NLO [Herrmann:2007ku, Harz:2023llw] exists, which provides the NLO (SUSY) QCD calculation to DM (co-)annihilation. The programs furthermore compute the direct detection cross sections. For further details and codes including indirect DM searches, we refer to [Arina:2020alg].
In this paper, we present a new tool, the C++ code RelExt. The code computes the relic density based on the freeze-out mechanism for arbitrary models featuring a DM candidate stabilized by a discrete $\mathbb{Z}_{2}$ symmetry. The user solely has to provide the FeynRules [Alloul:2013bka] model files, and the relic density is then automatically computed by RelExt. In this first release, the relic density is computed to leading-order in the thermally averaged cross section of the DM (co-)annihilation channels. The distinctly new feature of the code w.r.t. existing codes is the scan procedure. The code implements an efficient scan algorithm that searches for parameter configurations that generate the correct relic density within the experimentally allowed limits, provided the model under investigation is able to saturate the relic density. The code thereby provides a powerful tool to further constrain DM models by requiring the reproduction of the correct DM relic density. This can be used to subsequently investigate the phenomenological implications of this constraint for current and future collider and DM experiments on the one hand. On the other hand, this kind of investigation gives guidelines for further model building, in case the favored DM fails to reproduce the measured DM relic density.
@@
  • Having written the code ourselves, allows us to consistently combine it with the other codes generated by our collaboration to test new physics models with respect to particle, astroparticle and cosmological constraints, in particular our scanning tool ScannerS [Coimbra:2013qq, Muhlleitner:2020wwk], which checks for theoretical and experimental constraints of extended Higgs sector models, and the code BSMPT [Basler:2018cwe, Basler:2020nrq, Basler:2024aaf], which investigates BSM Higgs sectors with respect to their phase history in the early universe and possibly related gravitational waves signals.
At present, the following beyond-SM models featuring DM candidates have been pre-implemented and tested: the complex singlet extension of the Standard Model (CxSM) [Coimbra:2013qq, Costa:2015llh, Barger:2008jx, Gonderinger:2012rd, Muhlleitner:2017dkd, Chiang:2017nmu, Egle:2022wmq, Egle:2023pbm], the Next-to-2-Higgs-Doublet Model (N2HDM) in its dark doublet phase (DDP) [Engeln:2020fld, Azevedo:2021ylf], the model CP in the Dark (CPVDM) [Azevedo:2018fmj, Biermann:2022meg, Biermann:2023owb], the Two-Real-scalar-Singlet Model (TRSM) [Ghorbani:2014gka, Robens:2019kga], and the BDM5 [Huang:2020ris, Capucha:2022kwo]. New models, as stated above, can be included by providing the corresponding FeynRules model files. Note that the code at present does not handle models with more than one DM candidate. Furthermore, models where DM annihilation processes and the corresponding back-reactions differ (due to CP violation e.g.) are not supported either. Such extensions will be included in future releases.
@@ 2 Code Overview
This first part of the code allows the user to implement new models. This is done by providing the FeynRules [Alloul:2013bka] files for the model to be implemented. The code uses the packages FeynArts [Hahn:2000kx] and FeynCalc [Shtabovenko:2020gxv, Shtabovenko:2023idz] to automatically generate all tree-level DM annihilation amplitudes relevant for the calculation of the DM relic density. In this first release of the code these are the squared amplitudes for all tree-level DM $2\to 2$ (co-)annihilation processes. These amplitudes will be saved together with additional necessary information about the model in C++ files, which can then be compiled and used by the second part of the code. In case no new model is provided, the second part of the code can still be used with the already implemented models that are shipped with the code. The present version of the code is able to handle models with one $\mathbb{Z}_{2}$ symmetry, i.e. with one DM candidate.
@@
The second part of the code does not require Mathematica [Mathematica] and is purely written in C++. It performs the scans in the parameter spaces of the newly generated or already implemented models which search for parameter configurations that saturate the experimental limits on the relic density or that at least do not exceed it, in case weaker conditions are required. For this, the necessary thermally averaged cross sections are computed. This comprises also the computation of the needed total widths required for the virtual particle exchanges, and of the running couplings entering the interaction vertices of the scattering amplitudes. The code is designed such that it can be run with different settings and functions via a main.cpp file found within each model folder. The main.cpp allows the user to perform searches in the parameter space of a model using the provided functions (for more details see Sec. 3.4 and App. A). The values of the calculated relic densities and the corresponding relevant input parameters are saved in the output file specified by the user.
@@ 3.1 System Requirements
  • Mathematica v12[Mathematica] or higher is required for the implementation of new models. For the usage of implemented models, Mathematica and the packages listed here below are optional.
  • The Mathematica package FeynRules [Alloul:2013bka] is used to compute the Feynman rules for new models and to generate the required input files, which will be used by our code to implement new models.
  • FeynArts v3.11 [Hahn:2000kx] and FeynCalc v9.3.1 [Shtabovenko:2020gxv, Shtabovenko:2023idz] are used to calculate the squared amplitudes of the DM annihilation processes.
@@ 3.2 Download
  • base_Model
  • dataInput
  • dataOutput
  • include
    This directory contains the header files of the code.
@@
  • md_cpvdm
    This directory contains the already implemented model CPVDM [Azevedo:2018fmj, Biermann:2023owb].
  • md_cxsm
    This directory contains the already implemented model CxSM [Coimbra:2013qq, Costa:2015llh, Barger:2008jx, Gonderinger:2012rd, Muhlleitner:2017dkd, Chiang:2017nmu, Egle:2022wmq, Egle:2023pbm].
  • md_DDP
    This directory contains the already implemented model N2HDM in its dark doublet phase [Engeln:2020fld, Azevedo:2021ylf].
  • md_trsm
    This directory contains the already implemented model TRSM [Ghorbani:2014gka, Robens:2019kga].
  • md_bdm5
    This directory contains the already implemented model BDM5 [Huang:2020ris, Capucha:2022kwo].
  • numdata
    This directory contains the numerical data for the computation of the effective degrees of freedom [GONDOLO1991145].
  • sources
@@ 4 How to Implement a new Model
    The first input parameter is the path to the FeynRules directory. The second one is the name of the FeynRules model file, while the last input is the name of the Lagrangian variable, which is defined in the model file. This command will automatically search the model file in \$RelExt and generate all relevant (co-)annihilation amplitudes for DM freeze-out. The generated md_[ModelName] directory contains the following files and folders:
    • FR_modfiles
@@
    The last step is to call in $\texttt{\textdollar RelExt}/\texttt{build}$
@@ 5 Structure and Description of the Algorithms
The BSM models that provide DM candidates depend on a set of input parameters that can be more or less lengthy. Depending on the model and the interplay of the parameter values and the constraints on the models, more or less extensive parameter scans have to be performed, in order to find viable relic density values and to perform meaningful phenomenological investigations. A main objective of the code is therefore to solve the Boltzmann equation given in terms of the yield $Y$ by [PhysRevD.56.1879]
@@
Here, $\lambda$ is the Källén function, $\theta$ the azimuthal scattering angle, and $\mathcal{M}_{ij\rightarrow kl}$ the $2\rightarrow 2$ (co-)annihilation matrix element. After solving Eq. (5.1) from $x=0$ to $x=x_{0}$, the relic density can be calculated using [BELANGER2007367]
@@ 5.1 Amplitude Generation
where $m_{j}$ and $\Gamma_{j}$ are the mass and total decay width, respectively, of the mediator particle $j$. The code automatically identifies all $s$-channel mediator particles and calculates their total decay widths. We will give an overview on how the decay widths are obtained in Sec. 5.2. Additionally, in some kinematic regions it is possible to encounter $t/u$-channel divergencies. To regulate these divergencies we use a similar approach to MicrOMEGAs and introduce a width via $\Gamma_{j}=m_{j}/1000$ [Alguero:2023zol]. Using this width leads to a deviation of less than 1% outside of the divergent regions compared to the zero width result.
@@ 5.2 Computation of the Total Widths
In the computation of the total widths for scalar and pseudoscalar mediator particles, we include the available dominant higher-order corrections to the decay widths, as explained in more detail in the following. First of all, for the scalar/pseudoscalar mediator particles, analytical formulae are implemented in RelExt and used for the calculation of the partial decay widths, which are summed up to the total widths. We follow here mostly the implementation of HDECAY [Djouadi:1997yw, Djouadi:2018xqq], a code for the computation of the Higgs decays for the SM and the minimal supersymmetric SM extension (MSSM), including state-of-the-art higher-order corrections, and take the following formulae given in Ref. [Spira:2016ztx], unless stated otherwise. Here, we include the coupling modification factors of the scalar/pseudoscalar mediator to the SM particles according to the model under investigation. For the decays into lepton final states, we take the tree-level formula given in Eq. (21). Since the QCD interactions are not affected by modifications of the Higgs sector, we include higher-order QCD corrections in the decays into quarks. We interpolate here between the large Higgs mass region and the threshold region, and we include the relative QCD corrections $\delta_{\text{QCD}}$ and the relative top-quark induced contributions $\delta_{t}$, as given in Eq. (23). For the large Higgs mass region, we apply formula Eq. (15) for $\delta_{\text{QCD}}$, using running $\overline{\mbox{MS}}$ quark masses both at tree-level and at higher order. These and the running strong coupling constant $\alpha_{s}$ with five active flavours are taken at the scale of the mass of the decaying Higgs, thereby absorbing large logarithms. The $\delta_{t}$ corrections differ for the scalar and pseudoscalar case and are given after Eq. (23). In the threshold region, finite quark mass effects become relevant, and we apply Eq. (25). The QCD corrections are given in Eq. (17) for the scalar and in Eq. (26) for the pseudoscalar case, where we now take quark pole masses everywhere, both at leading and at higher order. The coupling modification factor for the Higgs coupling to the quarks in Eq. (23) and Eq. (25), respectively, is automatically replaced by the corresponding one for the model under consideration.
@@
Additionally, off-shell decays into massive gauge bosons $V$ ($V=W^{\pm},Z$) are taken into account. In a CP-conserving theory, they are only allowed for the scalar mediators. For energies $\sqrt{s}\geq 2m_{V}$, we apply the on-shell decay formula, given by Eqs. (34–36). Since the QCD corrections are of moderate size, we do not include them in our calculation. This also ensures a smooth transition to the off-shell decays, where no QCD corrections are available in a compact form. For energies below the production of any on-shell gauge bosons, $\sqrt{s}\leq m_{V}$, we apply Eqs. (37–39) for the production of two off-shell gauge bosons. In the energy range $m_{V}\leq\sqrt{s}\leq 2m_{V}$, we compute the decay into one off-shell and one on-shell gauge boson, using Eqs. (51,52), given in [Spira:1997dg].
@@
In BSM models, further Higgs decays are possible, which, depending on the parameter region, can become sizeable, or even dominant and therefore have to be included in the total widths as well. These can be e.g. decays into a lighter Higgs plus gauge bosons, lighter Higgs pairs, new heavy fermions or new heavy gauge boson pairs. The code computes all $1\rightarrow 2$ non-SM-like decays and adds them to the decay widths into SM particles to obtain the total width.
@@ 5.3 Computation of the Thermally Averaged Cross Section
With the generated amplitudes we move on to the computation of the TAC given in Eq. (5.4). We start by integrating Eq. (5.6) over $\cos\theta$. For this, we compute a first estimate of the integral using a composite Simpson’s $3/8$ rule [10.5555/1403886]. We use this estimate as part of the termination criterion for the adaptive Simpson’s $3/8$ method with which we compute the cross section. The use of the estimate in the termination criterion ensures a rapid convergence in areas where the integrand is small. Next, we need to integrate the numerator of Eq. (5.4). In order to do this we first sort the (co-)annihilation processes by the magnitude of their rest masses via
@@
This allows us to adjust the lower integration bound accordingly for each process. Additionally, we neglect channels which are heavily Boltzmann suppressed, i.e., that do not fulfill the criterion [Belanger:2006is]
@@
with the recommended value $B_{\epsilon}=10^{-6}$. Further, we approximate the modified Bessel functions of the second kind via [zbMATH03273551] [^8] To ensure that the precision of our results is not influenced by this, we use this approximation only for $x>5$.
[…]
This approximation results in an increase of the numerical stability of the TAC for large values of $x$ and decreases the computational effort. Similar to the $\cos\theta$ integration, we compute a first estimate of the TAC integral using the Kronrod 61 point method [10.5555/1403886] for the regions outside of $s$-channel resonance peaks. As in the $\cos\theta$ integration, we use this estimate in the termination criterion to compute the integral using an adaptive Gauss-Kronrod 15 point method [10.5555/1403886]. Since resonance peaks are challenging to deal with numerically, we integrate them separately if they are close enough to the energy of the rest mass of the considered channel, i.e. if the condition